# Robust Polyp Detection and Diagnosis through Compositional Prompt-Guided Diffusion Models

### v1

@@ 4.2.2 Polyp Classification
Figure 4: Radar chart illustrating the performance comparison between ResNet models trained on the original imbalanced dataset and the augmented balanced dataset.
@@
Figure 6: Left: Images generated with masks alone, masks with a text description (“polyp, reddish color”), and real images. Right: Images generated with bounding boxes (Bbox) alone and Bbox with text descriptions: Text1 (“polyp, 0-Isp”) and Text2 (“polyp, the surface is congested and granular”).
In the original imbalanced dataset experiment, we evaluated the performance of a ResNet classification model trained on this imbalanced dataset. The class imbalance negatively impacted the model’s performance, resulting in lower accuracy, especially in identifying malignant polyps. To address this issue, we conducted an experiment using an augmented balanced dataset. Additional images of malignant polyps were generated using the prompt “a V* polyp” creating a more balanced class distribution and mitigating the adverse effects of the original imbalance.
@@ 4.2.4 Qualitative Diversity Evaluation
Fig. 6 visually evaluates the diversity of generated samples, showcasing distinct textures, shapes, and sizes of polyps created using compositional prompts. Fig. 7 illustrates the transition during denoising, where the segmentation mask initially guides low-frequency structures, such as shape and position, while text descriptions refine high-frequency details like texture and pathology. This transition is controlled by a weighting scheme, with the text prompt weight $\lambda_{t}^{\text{text}}=(1+\cos(\pi\cdot t/T))/2$ increasing over time, and the mask prompt weight $\lambda_{t}^{\text{mask}}=1-\lambda_{t}^{\text{text}}$ ensuring a smooth shift between prompts.
@@ 5 Discussion
Recent advancements in generative models have shown promise for natural image synthesis [11]. However, challenges remain in medical imaging, particularly due to the complexity and variability of medical data [46]. Medical data are dynamic, evolving over time and varying across institutions, limiting the availability of comprehensive longitudinal datasets for training robust models. To address these challenges, our $\text{PSDM}^{\text{M}}$ strategy integrates multimodal data through compositional prompts, demonstrating substantial benefits, especially in complex datasets like PolypGen.
A limitation of our approach is the absence of established objective metrics for quantitatively evaluating the fidelity of synthetic medical images. While our model demonstrates the ability to generate visually high-quality images, a lack of rigorous quantitative evaluations, such as fidelity metrics or comparisons with clinical benchmarks, leaves the true clinical relevance of the generated images uncertain.
Looking forward, integrating synthetic data into clinical workflows presents a promising yet challenging frontier. Synthetic datasets offer opportunities to address class imbalances and simulate rare conditions. However, ensuring clinical interpretability, reliability, and utility for decision-making will be crucial for adoption in healthcare settings. Future research should focus on robust evaluation frameworks to assess the visual quality and clinical relevance of synthetic images.

### v2

@@ 4.2.2 Polyp Classification
Figure 4: Radar chart illustrating the performance comparison between ResNet models trained on the original imbalanced dataset and the augmented balanced dataset, with exact metric values overlaid.
@@
Figure 6: Left: Images generated with masks alone and masks with a text description (“polyp, reddish color”). Right: Images generated with bounding boxes (Bbox) alone and Bbox with text descriptions (“polyp, the surface is congested and granular”). The top-left inset highlights the lesion region.
In the original imbalanced dataset experiment, we evaluated the performance of a ResNet classification model trained on this imbalanced dataset. The class imbalance negatively impacted the model’s performance, resulting in lower accuracy, especially in identifying malignant polyps. Moreover, simply replicating minority-class samples can lead to overfitting on duplicated instances, further degrading generalization [49]. To address this issue, we conducted an experiment using an augmented balanced dataset. Additional images of malignant polyps were generated using the prompt “a V* polyp” creating a more balanced class distribution and mitigating the adverse effects of the original imbalance.
@@ 4.2.3 Polyp Detection
Table 6: Quantitative Evaluation Metrics
Method Metrics
FID ↓ SSIM ↑
ArSDM 200.35 0.4489
Mask_C 175.93 0.6115
Pix2pix 271.44 0.5759
$\text{PSDM}^{\text{S}}$ (Ours) 120.64 0.5770
$\text{PSDM}^{\text{M}}$ (Ours) 111.52 0.5839
@@ 4.2.4 Qualitative Diversity Evaluation
Fig. 6 visually evaluates the diversity of generated samples, showcasing distinct textures, shapes, and sizes of polyps created using compositional prompts. The enlarged lesion regions highlight that by introducing semantic control through textual prompts, we can generate more diverse samples. Fig. 7 illustrates the transition during denoising, where the segmentation mask initially guides low-frequency structures, such as shape and position, while text descriptions refine high-frequency details like texture and pathology. This transition is controlled by a weighting scheme, with the text prompt weight $\lambda_{t}^{\text{text}}=(1+\cos(\pi\cdot t/T))/2$ increasing over time, and the mask prompt weight $\lambda_{t}^{\text{mask}}=1-\lambda_{t}^{\text{text}}$ ensuring a smooth shift between prompts.
4.2.5 Quality Assessment of Generated Samples
Table 6 compares the quality of generated samples using FID and SSIM metrics. Our $\text{PSDM}^{\text{M}}$ achieved the lowest FID score (111.52), clearly outperforming other methods such as ArSDM (200.35) and Pix2pix (271.44), and moderately better than our own $\text{PSDM}^{\text{S}}$ (120.64). In terms of SSIM, $\text{PSDM}^{\text{M}}$ (0.5839) was second-best, close to Mask_C (0.6115), and notably higher than ArSDM (0.4489). These findings confirm that our approach provides good balance between visual fidelity and structural accuracy.
To further assess the clinical realism of our PSDM-generated images, we performed a blinded evaluation in which an experienced gastroenterologist (15 years of endoscopy practice) rated each image’s realism on a 5-point Likert scale. We randomly selected 150 images from each of two sets: Segmentation Mask Only set ($\text{PSDM}^{\text{S}}$) and Segmentation Mask-Text Description set ($\text{PSDM}^{\text{M}}$). Both sets yielded mean realism scores above 4, with $\text{PSDM}^{\text{S}}$ achieving a mean of 4.347 and $\text{PSDM}^{\text{M}}$ a mean of 4.447.
4.3 Ablation Studies on Prompt Integration and Weighting Schemes
To validate our coarse‐to‐fine compositional prompt strategy and assess the influence of different $\lambda$‐schedules (Eq. 3), we conducted two ablation studies: (i) measuring the efficacy of transitioning from coarse to fine prompts, and (ii) quantifying the impact of $\lambda$‐schedules on LPIPS and 1-SSIM. Because prompts at the same granularity are alternatives—segmentation masks and bounding boxes both encode spatial layout, and prototype images are needed only when textual descriptions are unavailable. We therefore examine two non‐overlapping combinations (mask + text and bbox + vague class) to highlight how integrating coarse localization with fine semantic cues enhances generation quality.
4.3.1 Effectiveness of Coarse-to-Fine Prompt Integration
We examine two coarse-to-fine transitions—segmentation mask $\rightarrow$ text prompts and bounding box $\rightarrow$ vague class prompts. Table 7 compares text-only, mask-only, and several mask $\rightarrow$ text schedules. Text-only guidance yields the highest semantic diversity (LPIPS = 0.6294) but the lowest structural coherence (1–SSIM = 0.8836), whereas mask-only guidance ensures structural fidelity (1–SSIM = 0.8867) at the expense of diversity (LPIPS = 0.5815). Among the mixed schemes, the linear (mask $\rightarrow$ text) schedule—which linearly decreases mask weight from 1 to 0 while increasing text weight from 0 to 1 over 200 diffusion steps—best balances both metrics (LPIPS = 0.5953, 1–SSIM = 0.9260). In contrast, the reverse linear (text $\rightarrow$ mask) transition underperforms (LPIPS = 0.5893, 1–SSIM = 0.8920) because later mask enforcement overrides earlier semantic cues. Likewise, Table 8 shows that vague class-only prompts produce poor diversity and coherence (LPIPS = 0.3887, 1–SSIM = 0.6085), while bbox-only prompts secure strong structure (1–SSIM = 0.8823) but lack semantic richness (LPIPS = 0.5914). All combined bbox $\rightarrow$ vague class schedules outperform single-cue baselines, with linear (bbox $\rightarrow$ vague class) achieving the best trade-off (LPIPS = 0.6294, 1–SSIM = 0.8866) by preserving spatial layout initially and gradually introducing semantic detail. These results rigorously confirm that a coarse-to-fine progression—first applying structural guidance (mask or bbox) then semantic refinement (text or vague class)—optimally enhances both the structural coherence and semantic diversity of generated polyp images.
Table 7: Impact of segmentation mask and textual prompts on image generation. Best results are bold, second-best are underlined.
Text Mask $\lambda$ Scheme LPIPS ($\uparrow$) 1–SSIM ($\uparrow$)
Single 0.6294±0.0687 0.8836±0.1090
0.5815±0.0743 0.8867±0.1135
Compositional stage (mask→text) 0.5922±0.0708 0.8884±0.1150
cosine (mask→text) 0.5947±0.0645 0.8969±0.1054
linear (mask→text) 0.5953±0.0655 0.9260±0.0860
linear (text→mask) 0.5893±0.0679 0.8920±0.1121
Table 8: Impact of bounding box and vague class prompts on image generation. Best results are bold, second-best are underlined.
Vague Class Bbox $\lambda$ Scheme LPIPS ($\uparrow$) 1–SSIM ($\uparrow$)
Single 0.3887±0.0280 0.6085±0.0611
0.5914±0.0833 0.8823±0.0843
Compositional stage (bbox→class) 0.6037±0.0739 0.8827±0.0983
cosine (bbox→class) 0.6146±0.0469 0.8857±0.0644
linear (bbox→class) 0.6294±0.0687 0.8866±0.1090
linear (class→bbox) 0.6013±0.0366 0.8822±0.0525
4.3.2 Impact of Weighting Hyperparameters
[…]
Figure 8: Examples of unsuccessful synthetic polyp images generated by our model.
@@ 5 Discussion
Recent advancements in generative models have shown promise for natural image synthesis [11]. However, challenges remain in medical imaging, particularly due to the complexity and variability of medical data [50]. Medical data are dynamic, evolving over time and varying across institutions, limiting the availability of comprehensive longitudinal datasets for training robust models. To address these challenges, our $\text{PSDM}^{\text{M}}$ strategy integrates multimodal data through compositional prompts, demonstrating substantial benefits, especially in complex datasets like PolypGen. The robustness of our approach is further evidenced by cross-LLM comparisons. Specifically, when substituting GPT-4o with DeepSeek-V2 to extract key attributes from text using a structured, template-based prompting strategy, both models consistently achieved error-free performance. Our PSDM approach also demonstrates broad applicability beyond polyp detection and diagnosis, as it imposes no restrictions on image modality—such as CT, MRI, histopathology, or endoscopy—provided multi-modal priors (e.g., masks, boxes, reports) are available.
A limitation of our approach is a slight decrease in training‐set performance—for example, the Polyp‐CASCADE model’s mDice on Kvasir declines from 93.09% to 92.01%, as shown in Table 3.This decrease highlights that, although our synthetic augmentation markedly strengthens robustness to unseen PolypGen samples, it can slightly impair performance on data drawn from the original training distribution. The baseline model retains higher in‐distribution accuracy by relying solely on familiar examples but fails to generalize to clinically diverse scenarios. Besides, we illustrate some failure samples of our method in Fig. 8, which demonstrate that underrepresented or weakly supported prompts (e.g., “yellow polyp” or “blue polyp”) often lead to indiscriminate colorization of the entire scene or outright generation failure, reflecting a weak prior for rare color–lesion combinations and an inability to disentangle the foreground lesion from the mucosal background when co‐occurrence statistics are scarce; notably, more explicit, compound prompts such as “red polyp in yellow rectum” provide stronger conditional cues that partially restore correct localization and contrast.
Looking forward, integrating synthetic data into clinical workflows presents a promising yet challenging frontier. Synthetic datasets offer opportunities to address class imbalances and simulate rare conditions. However, ensuring clinical interpretability, reliability, and utility for decision-making will be crucial for adoption in healthcare settings. Future research should focus on robust evaluation frameworks to assess the visual quality and clinical relevance of synthetic images, as well as dynamic strategies to balance domain-specific performance with cross-institutional generalization.